OpenStudy (anonymous):

ok i found the tangent line which is y=5x-19...now i need to find f(6) and F'(6)..can someone help me please?...o and tha tangent line at (6,3) passes through the point (0,2)....

OpenStudy (anonymous):

i don't understand wat question ur asking...if u hav the actual ques, just copy paste it...i'll help u out :)

OpenStudy (anonymous):

ok here it is.....Find an equation of the tangent line to the graph of y = g(x) at x = 3 if g(3) = −4 and g'(3) = 5. which i got it to be y=5x-19 If the tangent line to y = f (x) at (6, 3) passes through the point (0, 2), find f(6) and f '(6).

OpenStudy (anonymous):

ok, for this u don't need any information from g(x) or g'(x) f'(6) is the slope of the line joining (0,2) and (6,3) , which is 1/6; now find the equation of this tangent line, and f(6) is the y value in the tangent equation . got it huh?

OpenStudy (anonymous):

yea so the reverse would be 6 right?

OpenStudy (anonymous):

so that would be \[y=5(6)-19?\]